OpenStudy (anonymous):

Simplify: (n + 3)^2 Can someone explain how I simplify this?

OpenStudy (anonymous):

\[(n+3)^2=(n+3)(n+3)=n(n+3)+3(n+3).\] Does that help?

OpenStudy (anonymous):

Very little. I thought it would've been either N^2+ 9 or N2+6?

OpenStudy (anonymous):

Not really! Let me take if from where I stopped \[=n^2+3n+3n+9=n^2+6n+9.\]

OpenStudy (anonymous):

In a similar manner, you can find a general formula for \((x+y)^2\) as \[(x+y)^2=x^2+2xy+y^2.\]

OpenStudy (anonymous):

in that sense, I think it would then be n2 + 9n + 9? I think... I'm not very good with letters unfortunately. So I could've done that wrong

OpenStudy (anonymous):

It would be \(n^2+2(3)n+3^2=n^2+6n+9.\)